• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 16:05:31 on November 05, 2008

    Entry number 246929

    keyword=Preliminary Left HRS Detector Position Measurements

    Last week I made some measurements of the detector z positions with
    respect to the S1 plane. Using S1 is convenient since this detector's
    location with respect to the VDC is fixed and unchanging.
    In reality, I could not measure each detector with respect to S1 individually
    so instead I measured various distances along the detector
    stack including the thickness of A1 and the gas Cerenkov.
    I did not remeasure the RICH, since Youcai did this earlier. The
    detectors were centered last week and all y (transverse direction)
    distances were zeroed. The following were measured from the detector's front plane to their back planes:
    Thickness of A1:           29.2 cm (11.5")
    Thickness of Gas Cerenkov: 90.2 cm (35.5")
    The gas Cerenkov was difficult to measure due to it's irregular shape. I am
    sure I can check my numbers against earlier measurements.
    Detector Location        Position relative to S1 (m)      Z position (m)
    S1 position                             0                    1.2873
    A1_front                           0.0762                    1.3635
    A1_back                            0.3683                    1.6556
    GC_front                           0.4318                    1.7191
    GC_back                            1.3335                    2.6208
    S2m front                          1.8542                    3.1415
    I was only able to measure the distance from S2m frame to the GC_back plane.
    I have some notes from Rob Feuerbach that there is a 1"
    distance between the active region of S2m above the frame position.

    Using 1.8542 m distance between S1 and S2m, the time-of-flight between
    the two detectors is about 6.2 ns.

    I will update the detector databases with these new numbers.

    A copy of this log entry has been emailed to: jiang,ycwang,chiran,brads